# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v3

@@ 4.1 Experimental Setup
We select 26 widely-used and high-performing LLMs. These models are categorized into three types based on their accessibility and purpose. The details about the implementation of models can be found in Appendix A2.
  • Proprietary LLMs: This category generally represents state-of-the-art. Included models are OpenAI’s GPT series [29], specifically OpenAI’s GPT-4o, GPT-4o-mini, GPT-4-Turbo, GPT-3.5-Turbo, Anthropic’s Claude-3.5-Sonnet, Claude3-Sonnet [1], Google’s Gemini1.5-Pro [32], and Alibaba’s Qwen-Max [43].
  • Open-Source General-Purpose LLMs: These LLMs excel in general domains and serve as a foundation for further research into scientific LLMs. We selected seven LLMs from diverse sources, ranging in size from 7B to 72B, including Qwen2-7/72B-Inst [43], Llama3-70B-Inst [8], Qwen1.5-7/14B-Chat [2], Llama3-8B-Inst [8], Llama2-13B-Chat [37], ChatGLM3-6B [7], Gemma1.1-7B-Inst, and Mistral-7B-Inst [36].
@@
  • Open-Source Scientific LLMs: These models have acquired specialized knowledge by training on scientific domain data. We selected models focused on biology and chemistry, including ChemDFM-13B [47], Galactica-6.7B/30B [35], ChemLLM-7B/20B-Chat [46], MolInst-Llama3-8B [12], LlaSMol-Mistral-7B [44], and SciGLM-6B [45].
Evaluation Mode.
In our experiments, the input begins with a system prompt that describes the types and categories of questions. We employ two evaluation settings: zero-shot and few-shot. The zero-shot setting evaluates models’ problem-solving capabilities without any prior examples, testing their ability to solve problems based on their own inherent knowledge. In the few-shot setting, models are provided with a limited number of examples before the test example to assess their capability to acquire and incorporate new information into their problem-solving processes.
Evaluation Criteria.
@@
We adopt diverse evaluation metrics, tailoring our assessment to different task types. When evaluating True/False, classification and multiple-choice questions, we use accuracy as the performance metric. For relation extraction questions, we use the $F_{1}$-score that combines precision and recall. For generative questions, we adopted different evaluation methods tailored to the characteristics of each task. Specifically, following the existing works [9, 44, 14, 12, 10], we calculate the average scores using BLEU [31] and ROUGE [24] metrics for the tasks of molecule captioning, protein captioning, and single cell analysis. For molecule generation, we compute the average of the exact match rate and fingerprint-based Tanimoto similarity [44]. For protein generation, we leverage the Smith-Waterman algorithm [33] to perform local sequence alignment of two protein sequences. For other generative tasks, we designed meticulous prompts for GPT-4o to evaluate the responses of LLMs. The scoring prompt templates are exhibited in Appendix A9. Considering the challenge of aggregating different metrics, we report the average rankings of LLMs in each task as the final score.
4.2 Results and Analyses
In this section, we report the performance of LLMs in the SciKnowEval dataset. Table 4 and Table A1.1 summarize the zero-shot performance rankings of LLMs at each level, offering valuable insights into the strengths and weaknesses exhibited by each model. We emphasize our key observations as follows, and the illustrative examples can be found in Appendix A10.
@@ 4.2 Results and Analyses
Table 4: Overall zero-shot performance of LLMs across five levels in four domains. A smaller value indicates a higher ranking. Bold results indicate the best results among all models, underline results indicate the second-best results, and blue results indicate the best results among the open-source models.
Categories Models L1 L2 L3 L4 L5 All Rank
Proprietary LLMs Claude3.5-Sonnet 2.70 4.80 4.00 2.90 2.27 3.71 1
GPT-4o 2.30 4.56 5.95 6.40 3.00 4.68 2
GPT-4-Turbo 6.40 6.12 8.59 7.50 5.09 6.88 4
Gemini1.5-Pro-latest 8.20 8.44 6.00 5.10 7.18 7.12 5
GPT-4o-mini 9.10 7.80 12.14 6.30 4.73 8.56 7
Qwen-Max 7.90 7.76 9.27 7.90 10.36 8.59 8
Claude3-Sonnet 9.20 8.92 10.82 9.60 6.00 9.17 9
GPT-3.5-Turbo 11.60 13.24 14.82 10.80 10.55 12.78 12
Open-Source General-Purpose LLMs Qwen2-72B-Inst 4.90 5.28 8.55 4.20 7.64 6.35 3
Llama3-70B-Inst 7.90 6.24 8.86 5.30 7.18 7.21 6
Qwen2-7B-Inst 12.40 11.40 14.14 9.70 14.09 12.46 10
Qwen1.5-14B-Chat 12.40 13.36 11.95 13.60 11.91 12.67 11
Llama3-8B-Inst 12.80 12.32 14.73 11.80 17.00 13.65 13
Qwen1.5-7B-Chat 15.70 15.60 17.50 16.60 17.82 16.59 17
Gemma1.1-7B-Inst 18.90 20.40 15.59 17.70 15.64 17.83 18
Mistral-7B-Inst 20.20 16.88 18.59 15.30 18.64 17.83 18
ChatGLM3-6B 19.00 20.56 18.64 18.00 17.64 19.08 20
Llama2-13B-Chat 21.80 18.56 21.73 18.00 17.45 19.64 22
Open-Source Scientific LLMs ChemDFM-13B 12.50 15.24 14.45 15.10 16.09 14.77 14
ChemLLM-20B-Chat 15.00 12.80 16.27 19.60 16.82 15.50 15
MolInst-Llama3-8B 17.40 15.88 12.41 16.80 18.73 15.62 16
Galactica-30B 17.20 21.48 16.09 22.80 19.82 19.35 21
SciGLM-6B 21.70 20.32 19.41 22.40 21.55 20.68 23
ChemLLM-7B-Chat 20.30 21.16 20.36 21.90 20.09 20.77 24
Galactica-6.7B 21.60 23.60 17.18 22.50 24.00 21.45 25
LlaSMol-Mistral-7B 22.60 23.84 20.59 25.90 20.91 22.62 26
Overall Performance.
@@
Proprietary LLMs, such as the GPT-4 series and Claude-3.5, have consistently demonstrated superior performance across these four domains, securing their highest overall rankings. Notably, Claude3.5-Sonnet has exhibited exceptional capability and adaptability in scientific domains. Open-source LLMs with larger scales, including Llama3-70B and Qwen2-72B, also exhibited comparable performance. In contrast, scientific LLMs performed moderately and only showcased strengths in a few tasks.
Performance on Each Level.
@@
L1 reflects the model’s memory of scientific knowledge. Proprietary LLMs, such as GPT-4o demonstrated the best capabilities in four domains, showcasing its extensive knowledge coverage. In the medium-scale ($\sim$10B) LLMs, ChemDFM emerged as one of the top open-source models by continuing pre-training and fine-tuning on a vast corpus of scientific literature. However, many scientific LLMs, such as LlaSMol-Mistral-7B, lagged behind, possibly due to overfitting caused by specific instruction fine-tuning.
@@
L2 measures the model’s comprehension ability within scientific contexts. GPT-4o and other proprietary LLMs showcased strong text comprehension performance, which also included open-source models like Qwen2-series. However, they struggled with tasks involving relation extraction.
L3 evaluates the model’s reasoning and computational abilities for scientific questions. In the biological domain, despite GPT-4o and Gemini1.5-Pro demonstrating relatively higher average rankings, they did not exhibit significant superiority in protein function prediction tasks. In the chemical domain, GPT-4o performed relatively better in tasks such as reaction prediction, retrosynthesis, and chemical calculation, but there remains substantial room for improvement in other tasks. Overall, all evaluated models need further enhancement in scientific computation.
@@
L4 highlights the model’s awareness of scientific safety. For harmful QA tasks in biology and chemistry, LLMs are expected to refuse to answer harmful scientific questions. Gemini1.5-Pro showed strong safety judgment, with refusal rates of 81.9% in chemistry and 100% in biology. Claude3.5-Sonnet showed similar performance. However, other models, including the GPT-4 series, underperformed. In molecular toxicity prediction, only a few LLMs exceeded 40% accuracy, revealing their limitations in assessing molecular toxicity. Lastly, in laboratory safety tests, proprietary models like GPT-4o excelled, showing promise for safe lab operations.
@@
L5 reflects the creative abilities of LLMs in real-world scientific scenarios, determining their potential in experimental protocol design, drug design, and so on. For the protocol design tasks in both biology and chemistry, we prompt GPT-4o to rate results from 1 to 5. However, despite proprietary models like GPT-4o outperforming others, almost no model achieved an average score of 3 out of 5. This indicates that existing models are still unable to generate high-quality experimental protocols. In the molecular generation tasks, scientific LLMs such as LlaSMol-Mistral-7B and ChemDFM-13B significantly outperformed other models, which we attribute to their training data involving molecular description and generation. In the protein design tasks, none of the current LLMs delivered satisfactory results, with average normalized Swith-Waterman alignment scores approaching zero. Additionally, performance bottlenecks were also encountered in the single-cell analysis task. In summary, the creative capabilities of LLMs related to molecules, proteins and cells require further improvement.
Table 5: 3-shot performance of LLMs on each task in the biology domain. Bold indicates performance improvement compared to the zero-shot setting, while gray signifies no gain. Blue results indicate a large improvement (>10%). The 3-shot results of other domains are provided in Table A2.
Domain Task Name GPT-4o Claude3 Llama3-8B Qwen1.5-7B MolInst-8B ChemLLM-7B
0-shot 3-shot 0-shot 3-shot 0-shot 3-shot 0-shot 3-shot 0-shot 3-shot 0-shot 3-shot
Bio. Bio LiterQA (L1) 83.71 +0.45 76.44 -0.72 74.82 +2.15 72.06 +0.69 72.82 +2.30 61.65 +8.30
Prot Prop Iden. (L1) 34.20 +10.53 27.93 +4.47 25.60 +3.73 26.40 -1.8 20.53 +3.54 22.20 +2.73
Protein Cap. (L1) 0.122 +0.005 0.118 +0.002 0.112 -0.003 0.100 +0.028 0.009 +0.153 0.065 +0.049
D-D RE (L2) 17.41 +0.22 16.02 +0.86 16.75 +0.26 16.93 -0.53 14.86 -0.69 6.47 +3.27
Bio JI (L2) 95.46 -0.53 96.20 -0.32 89.12 +5.92 87.33 +2.22 92.71 5.39 26.08 +39.92
C-D RE (L2) 37.70 +10.59 35.85 +8.5 28.90 +24.09 21.27 +6.42 50.26 -5.42 13.34 +11.08
G-D RE (L2) 36.20 -1.15 29.18 -5.75 39.16 -0.52 8.74 +24.72 10.38 +0.75 12.65 +6.34
Bio DU (L2) 99.40 +0.00 98.67 -0.72 98.79 -0.23 97.58 -0.48 96.50 +0.84 96.01 -0.84
Bio HV (L2) 94.82 -0.16 94.98 +0.00 90.45 +0.16 85.60 +0.00 89.48 -5.18 61.97 +16.99
Bio RI (L2) 97.69 +0.00 96.60 +0.47 96.76 -0.43 95.22 -0.16 93.21 -0.46 91.67 -0.47
Solu. Pred (L3) 48.31 +5.31 46.86 +1.45 49.28 -0.49 49.28 +4.34 53.62 +0.00 46.86 +0.48
$\beta$-LA Pred (L3) 50.25 +13.30 43.84 +24.63 50.25 +5.42 27.59 +24.63 47.78 +20.20 61.58 -13.80
Fluo. Pred (L3) 51.72 +27.10 51.72 +28.58 49.75 +9.36 46.31 +14.28 44.83 +22.17 20.69 +33.00
GB1 Pred (L3) 32.54 -0.40 24.40 +4.31 19.14 +4.78 13.40 +15.79 28.23 -0.24 16.27 +15.79
Stab. Pred (L3) 24.02 +1.96 30.39 -3.43 24.51 +2.45 22.55 15.69 28.92 -0.59 26.47 +4.41
Prot-Prot Inter. (L3) 35.27 -6.77 27.54 -1.94 21.26 +11.59 22.22 +3.87 25.60 +6.28 20.77 10.15
Bio Cal. (L3) 58.33 -8.33 38.33 -3.33 33.33 +5.00 36.67 +10.00 41.67 -3.34 21.67 +15.00
Proteotox. (L4) 85.88 -0.59 40.88 +20.3 38.24 +3.13 39.71 +10.49 37.06 +0.98 22.06 +9.90
Bio Safe Test (L4) 86.67 +0.44 71.11 +1.57 64.44 +5.15 67.78 +3.87 65.56 +1.45 61.11 -5.44
Protein Des. (L5) 0.010 -0.003 0.008 -0.001 0.000 +0.001 0.000 +0.004 0.000 +0.004 0.000 +0.004
4.3 Discussions
SciKnowEval exhibits Sufficient Difficulty and Challenge.
@@ 4.3 Discussions
Firstly, our results indicate that in zero-shot setting, proprietary models consistently outperform other open-source models. Moreover, there is a noticeable positive correlation between model size and performance, e.g., Galactica-30B outperforms Galactica-6.7B, and Qwen1.5-14B-Chat exceeds the performance of Qwen1.5-7B. Secondly, by examining the detailed scores of GPT-4o across various tasks, it is evident that SciKnowEval spans multiple levels of difficulty. For most tasks at the L1 and L2 levels, GPT-4o achieves accuracies above 85%. However, GPT-4o struggles with tasks at the L3 and L5 levels, particularly those involving molecular SMILES and protein sequences. Lastly, our carefully designed L4 level, aimed at evaluating the safety of LLMs, introduces a novel challenge compared to other benchmarks such as SciEval and SciAssess. We observed that GPT-4o often failed to reject harmful questions in the Harmful QA task, presenting a potential risk of misuse.
Few-shot Setting enhances Model Performance across Tasks.
For the few-shot setting, we selected two competitive models from each of the three categories of LLMs, and evaluated them on most of the tasks in our datasets. Table 5 summarizes the task performance shifts of these LLMs under a 3-shot setting. For most tasks involving multiple-choice and true/false questions, we observed significant performance gains, especially for tasks at the L3 level. For instance, GPT-4o and Claude3-Sonnet saw accuracy increases of 27.10% and 28.58%, respectively, in the fluorescence prediction task. This indicates that few-shot settings can substantially enhance models’ scientific reasoning and computational abilities.
Incremental Pre-training or Fine-tuning on Scientific Corpus show Promise. We compared two pairs of models: 1) Llama2-13B vs. ChemDFM-13B, and 2) Mistral-7B-Inst vs. LlaSMol-Mistral-7B. We can observe that ChemDFM-13B, built on the Llama-13B framework and further pre-trained and fine-tuned on a corpus of 34 billion tokens from scientific literature, significantly outperformed Llama2-13B-Chat. Similarly, fine-tuning with instruction data for molecule generation and description enabled LlaSMol-Mistral-7B to surpass Mistral-7B-Inst at the L1 and L5 levels in the chemical domain.
@@
Evaluating OpenAI o1 with SciKnowEval. Recently, OpenAI released a new LLM, o1 [28], which excels at complex task reasoning, especially in the fields of science, mathematics, and programming. In a series of challenging benchmarks, o1 delivered outstanding results and even surpassed human experts in PhD-level scientific Q&A sessions. To further assess o1’s capabilities in scientific domains, we used a subset of the SciKnowEval dataset to evaluate the performance of o1-preview and o1-mini. The subset consists of 1,775 challenging questions that GPT-4o-mini fails to answer correctly. The evaluation results are presented in Figure A1. Through analyzing the quantitative results and several cases, we have three key findings: (1) By generating hidden chain-of-thoughts (CoT) during model inference, o1 has a significant improvement in answering questions related to scientific computation and reasoning, however, it occasionally falls into the traps of reasoning problems, especially those involving complex physical principles and laws. (2) o1 has integrated safety rules into its CoT, improving the model’s ability to align with safety, but it still lacks sufficient knowledge regarding certain substances (such as uncommon toxic compounds and viruses) and may output harmful responses. (3) In addition to scientific reasoning and safety alignment, the improvement of o1 in scientific knowledge memory, knowledge understanding, and knowledge application is limited. See Appendix A11 for more evaluation results and case analyses.

### v4

@@ 3.1 Experimental Setup
We select 20 widely-used and high-performing LLMs. These models are categorized into three types based on their accessibility and purpose. The details about the implementation of models can be found in Appendix A3.
  • Proprietary LLMs: This group includes state-of-the-art LLMs developed by leading organizations. Specifically, we evaluate several models from OpenAI, including GPT-4o, GPT-4o-mini, and GPT-4.1 ouyang2022training , as well as more recent reasoning models such as o3-mini and o4-mini openai2024o3o4mini . From Anthropic, we include both Claude-4-Sonnet and its reasoning-mode variant Claude-4-Sonnet-thinking anthropic2024claude .
  • Open-Source General-Purpose LLMs: This category comprises LLMs that demonstrate strong performance in general domains and are commonly used as the foundation for developing scientific LLMs. In this study, we evaluate eight LLMs of varying scales (ranging from 7B to 671B parameters), sourced from multiple organizations. The selected models include Qwen2-7B-Instruct yang2024qwen2 , Qwen2.5-72B-Instruct yang2024qwen25 , QwQ-32B qwq32b , and Qwen3-8B-thinking yang2025qwen3 developed by Alibaba, Llama3-8B-Instruct dubey2024llama and Llama4-Scout meta2025llama from Meta, as well as DeepSeek-R1 guo2025deepseek and DeepSeek-V3 liu2024deepseek from DeepSeek.
@@
  • Open-Source Scientific LLMs: These models have acquired specialized knowledge by training on scientific domain data. In our evaluation, we focus on models tailored for the scientific domains covered by SciKnowEval, including ChemDFM-13B zhao2024chemdfm , ChemLLM-20B-Chat zhang2024chemllm , MolInst-Llama3-8B fang2023mol , LlaSMol-Mistral-7B yu2024llasmol , and SciGLM-6B zhang2024sciglm .
Evaluation Setting
In our experiments, the input begins with a system prompt describing the types and categories of questions. We then employ a zero-shot evaluation setting, where the model is presented only with the question itself and no additional examples, in order to assess its problem-solving capabilities based solely on its inherent knowledge.
Evaluation Criteria
@@
We adopt diverse evaluation metrics, tailoring our assessment to different task types. When evaluating True/False, classification and multiple-choice questions, we use accuracy as the performance metric. For relation extraction questions, we use the $F_{1}$-score that combines precision and recall. For generative questions, we designed meticulous prompts for GPT-4o to evaluate the responses of LLMs. The scoring prompt templates are exhibited in A9. We normalize the results of all evaluation metrics to the range of 0 to 1. We then compute the average score for each level, as well as the overall average score across all levels.
@@
Table 1: Overall zero-shot performance of LLMs across five levels in four domains. The scores are from 0 to 1, a higher score means better performance Bold results indicate the best results among all models, underline results indicate the second-best results, and blue results indicate the best results among the open-source models.
Categories Models L1 L2 L3 L4 L5 OverAll Rank
Proprietary
LLMs
o4-mini 0.859 0.843 0.589 0.768 0.491 0.710 1
o3-mini 0.860 0.839 0.597 0.697 0.486 0.696 2
GPT-4.1 0.863 0.844 0.525 0.694 0.472 0.679 3
GPT-4o 0.840 0.833 0.493 0.672 0.410 0.650 6
Claude-4-Sonnet-thinking 0.767 0.851 0.462 0.679 0.410 0.634 9
GPT-4o-mini 0.792 0.802 0.453 0.668 0.371 0.617 12
Claude-4-Sonnet 0.725 0.825 0.425 0.686 0.420 0.616 13
Open-Source
General-Purpose
LLMs
DeepSeek-V3 0.829 0.835 0.520 0.652 0.448 0.657 4
DeepSeek-R1 0.827 0.833 0.477 0.650 0.447 0.647 7
QwQ-32B 0.818 0.842 0.566 0.638 0.417 0.655 5
Qwen2.5-72B-Instruct 0.826 0.825 0.479 0.678 0.384 0.638 8
Qwen2-7B-Instruct 0.760 0.791 0.379 0.632 0.255 0.564 14
Llama4-Scout 0.817 0.791 0.541 0.638 0.379 0.633 10
Qwen3-8B-thinking 0.803 0.818 0.453 0.651 0.370 0.619 11
Llama3-8B-Instruct 0.756 0.580 0.361 0.677 0.019 0.479 17
Open-Source
Scientific
LLMs
ChemDFM-13B 0.717 0.759 0.388 0.566 0.174 0.521 15
ChemLLM-20B-Chat 0.711 0.734 0.354 0.515 0.092 0.481 16
MolInst-Llama3-8B 0.726 0.673 0.381 0.553 0.042 0.475 18
SciGLM-6B 0.622 0.634 0.283 0.423 0.035 0.399 19
LlaSMol-Mistral-7B 0.359 0.415 0.189 0.192 0.021 0.235 20
3.2 Main Results
In this section, we report the performance of LLMs in the SciKnowEval dataset. Table 1 and Table A2 summarize the zero-shot performance rankings of LLMs at each level, offering valuable insights into the strengths and weaknesses exhibited by each model. We emphasize our key observations as follows.
Overall Performance
@@ 3.2 Main Results
Proprietary LLMs, such as GPT-4.1 and the GPT o-series, consistently demonstrate superior performance across these four domains, securing their highest overall rankings. Notably, o4-mini exhibits exceptional capability and adaptability in scientific domains. Open-source LLMs with larger scales, including DeepSeek-V3, DeepSeek-R1, and QwQ-32B, also exhibit comparable performance. In contrast, scientific LLMs perform moderately and only showcase strengths in a few tasks. It is particularly noteworthy that large reasoning models, whether proprietary or open-source, achieve outstanding performance. This advantage primarily stems from their deliberative and scalable reasoning capabilities.
Performance on Each Level
@@
L1 reflects the model’s memory of scientific knowledge. Proprietary LLMs, such as GPT-4.1, demonstrate the best capabilities in four domains, showcasing their extensive knowledge coverage. For open-source LLMs, large-scale models such as DeepSeek-V3 and Qwen2.5-72B-Instruct significantly outperform smaller models like Qwen2-7B-Instruct and Llama3-8B-Instruct. This advantage is likely attributed to their larger parameter capacity, which enables more comprehensive knowledge retention and better generalization. However, many scientific LLMs, such as LlaSMol-Mistral-7B, lag behind, possibly due to overfitting caused by specific instruction fine-tuning. We also find that reasoning models do not show performance gains at this level, possibly due to hallucinations introduced by their complex reasoning processes.
@@
L2 measures the model’s comprehension ability within scientific contexts. Claude-4-Sonnet-thinking demonstrates strong text comprehension performance, a strength also observed in open-source models such as QwQ-32B. Additionally, proprietary models like GPT-4.1 and GPT-4o also achieved outstanding performance in material, physics, and chemistry tasks due to their powerful instruction-following capabilities. However, almost all LLMs struggled with tasks involving relation extraction, which reveals a distinct contrast with other tasks in L2, especially for the biological ones.
L3 evaluates the model’s reasoning and computational abilities for scientific questions. o3-mini and o4-mini, benefiting from large-scale reinforcement learning, achieve the best performance at this level and demonstrate strong analytical capabilities. Despite these reasoning models achieving relatively high evaluation results and rankings, they still struggle with certain tasks, such as the “Stability Prediction” task in the biological domain and the “Molecular Structure Prediction” task in the chemical domain. Overall, all evaluated LLMs need further enhancement in scientific computation.
@@
L4 highlights the model’s awareness of scientific safety. For harmful QA tasks across all four domains, LLMs are expected to refuse to answer harmful scientific questions. o4-mini shows strong safety judgment, with refusal rates of 86.9% in material, 86.8% in physics, and 100% in biology. We attribute this to deliberative alignment guan2024deliberative , a novel safety alignment approach. However, other models, including GPT-4.1, perform worse in this aspect. In molecular toxicity prediction, only a few LLMs exceed 60% accuracy, revealing their limitations in assessing molecular toxicity. Lastly, in laboratory safety tests, proprietary models like GPT-4.1 excel, showing promise for safe lab operations.
@@
L5 reflects the creative abilities of LLMs in real-world scientific scenarios, determining their potential in experimental protocol design, material synthesis, and so on. For the protocol design tasks in both biology and chemistry, we prompt GPT-4o to rate results from 1 to 5, then map them to the range of 0 to 1 to get the final score. However, despite proprietary models like o4-mini and o3-mini outperforming others, none of the models reaches an average score of 3 out of 5. This indicates that existing models are still unable to generate high-quality experimental protocols. Additionally, performance bottlenecks are also observed in the specified band gap generation task in material, as well as the problem-solving task in physics. In summary, the creative capabilities of LLMs require further improvement.
Table 2: Model Performance and Rankings Across Scientific Domains. The scores are from 0 to 1, a higher score means better performance Bold results indicate the best results among all models, underline results indicate the second-best results, and blue results indicate the best results among the open-source models.
Models Biology Chemistry Material Physics
Score Rank Score Rank Score Rank Score Rank
o4-mini 0.6268 1 0.6800 2 0.7364 1 0.8922 1
o3-mini 0.5955 3 0.6861 1 0.7210 2 0.8910 2
GPT-4.1 0.6075 2 0.6291 4 0.7043 4 0.8548 5
GPT-4o 0.5758 6 0.6221 6 0.6555 9 0.8369 9
Claude-4-Sonnet-thinking 0.5148 14 0.6274 5 0.6939 5 0.8391 8
GPT-4o-mini 0.5777 5 0.5759 13 0.6007 11 0.8046 12
Claude-4-Sonnet 0.5661 9 0.6025 9 0.5970 13 0.8031 13
DeepSeek-V3 0.5833 4 0.6163 7 0.6799 7 0.8552 4
DeepSeek-R1 0.5663 8 0.5766 12 0.6913 6 0.8537 6
QwQ-32B 0.5532 12 0.6388 3 0.7203 3 0.8731 3
Qwen2.5-72B-Instruct 0.5696 7 0.5934 11 0.6514 10 0.8457 7
Qwen2-7B-Instruct 0.5291 13 0.5316 14 0.5458 14 0.7565 14
Llama4-Scout 0.5590 10 0.5972 10 0.6718 8 0.8281 10
Qwen3-8B-thinking 0.5564 11 0.6026 8 0.5979 12 0.8211 11
Llama3-8B-Instruct 0.4672 16 0.4727 18 0.4407 18 0.5369 18
ChemDFM-13B 0.5084 15 0.5314 15 0.5321 15 0.6129 15
ChemLLM-20B-Chat 0.4606 17 0.5071 16 0.4681 17 0.6104 16
MolInst-Llama3-8B 0.4592 18 0.4739 17 0.4718 16 0.6045 17
SciGLM-6B 0.4148 19 0.4007 19 0.3584 19 0.5344 19
LlaSMol-Mistral-7B 0.2133 20 0.2774 20 0.2398 20 0.3198 20
Performance across Domains
Table 2 shows the performance of the LLMs in each domain. In biology, materials, and physics, o4-mini consistently outperforms other LLMs, while o3-mini achieves the best performance in the chemistry domain. Most LLMs exhibit similar ranking trends across the four domains, reflecting the strong generalization ability of general-purpose language models. However, there are some exceptions. For instance, Claude4-Sonnet-thinking and QwQ-32B perform relatively poorly in biology, and DeepSeek-R1 shows inferior results in chemistry. Furthermore, GPT-4o-mini demonstrates comparatively stronger performance in biology. Finally, scientific LLMs that are fine-tuned on specific scientific tasks, such as ChemLLM-20B-Chat, do not show clear advantages in their corresponding domains. This may be attributed to outdated model versions and overfitting, highlighting important considerations for training domain-specific models.
3.3 Findings
SciKnowEval exhibits Sufficient Difficulty and Challenge
@@ 3.3 Findings
Firstly, our results indicate that in zero-shot setting, proprietary models consistently outperform other open-source models. Moreover, there is a noticeable positive correlation between model size and performance. Secondly, by examining the detailed scores of GPT-4o across various tasks, it is evident that SciKnowEval spans multiple levels of difficulty. For most tasks at the L1 and L2 levels, GPT-4o achieves accuracies above 85%. However, GPT-4o struggles with tasks at the L3 and L5 levels, particularly those involving molecular SMILES and protein sequences. Lastly, our carefully designed L4 level, aimed at evaluating the safety of LLMs, introduces a novel challenge compared to other benchmarks such as SciEval and SciAssess. We observe that GPT-4o often fails to reject harmful questions in the Harmful QA task, presenting a potential risk of misuse.
Incremental Pre-training or Fine-tuning on Scientific Corpus shows Promise
We compare the pair of models: Llama3-8B-Instruct vs. MolInst-Llama3-8B. We observe that MolInst-Llama3-8B, built on Llama3-8B-Instruct and further fine-tuned on Mol-Instructions fang2023mol , has a clear advantage at the biological and chemical tasks in L4 and L5, and text summary tasks in the biology and chemistry domains. It also has a clear advantage in some of the L3 tasks, like protein-protein interaction and valence electron difference calculation. In summary, compared to Llama3-8B-Instruct, MolInst-Llama3-8B shows better performance at most of the application tasks and molecular tasks.
Large Reasoning Models Exhibit Strong Scientific Reasoning and Safety Capabilities
@@
Recently, advanced large reasoning models (LRMs) such as the GPT o-series openai2024o3o4mini , DeepSeek-R1 guo2025deepseek , and DeepSeek-V3 liu2024deepseek are released, excelling in complex task reasoning, particularly in the fields of science, mathematics, and programming. In a series of challenging benchmarks, LRMs deliver outstanding results and even surpass human experts in PhD-level scientific Q&A sessions. O3-mini and O4-mini show leading performance in almost all aspects. Through analyzing the quantitative results and several cases, we have three key findings: (1) By generating hidden chain-of-thoughts (CoT) during inference, LRM shows a significant improvement in answering questions related to scientific computation and reasoning, though it occasionally falls into reasoning traps, especially with complex physical principles and laws. (2) LRMs integrate safety rules into their CoT, improving safety ability, but still lack sufficient knowledge regarding certain substances (e.g., rare toxic compounds, viruses), leading to harmful outputs. (3) Despite advances in reasoning and safety, improvements in scientific knowledge memory, understanding, and application remain limited.